Operations Report – December 17th

Name of person filing report: Shannon Rupert

Reason for Report: Routine/crew changeover

Non-nominal systems: Hab toilet has something wrong and it is not emptying correctly.

Action taken for non-nominal systems: We will be investigating tomorrow

Generator: Running from about 8 pm to 8 am. We looked into some of the issues surrounding the autostart last weekend and as usual, the power system decided to seek revenge. The router is showing as SOC connect for the generator but the online data is only showing VDC. Sooooo…. We need to figure out which parts of the system need reprogrammed.

ScienceDome Dual Split: on at 65 degrees overnight

Solar— Nominal (see generator notes for some issues with the power system as a whole)

Solar— SOC Last 24 hours:



Avg not showing the way it is currently operating

Propane Reading, station tank – filled Friday %

Propane Reading, director tank— filled Friday %

Propane Reading, intern tank— Filled Friday %

Propane Reading, generator— Filled Friday %

Ethanol Free Gasoline – 0 gallons

Water (static tank) – 550 gallons

Water in GreenHab – 200 gallons

Water in ScienceDome: 0 gallons

Water (Outpost tank) – 550 gallons

Hab toilet tank emptied: Yes, it has been emptied every two days this week as there is a blockage.

Sojourner rover used: no

Hours: not noted

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Notes on rovers: nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): none

Reason for use: n/a

Oil Added? no

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used today: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, to town and on SWB field trip to Salt Wash

CrewCar used and why, where? No. It has been in Grand Junction.

Luna used and why, where? Yes, to town

General notes and comments: Camera with alarm (to alert us to trespassers) was installed on the side of the Hab but was too far away to trigger the alarm so we are moving it soon to the old HughesNet mount near the GreenHab

Summary of internet: Nominal.

Summary of suits and radios: See crew operations reports.

Campus wide inspection, if action taken, what and why? Nothing to report

Summary of general operations: Conduit work delayed

Summary of Hab operations: As noted above the toilet has a problem that we need to investigate. I ordered a new humidifier for the upper deck. The one we had was not operational nor was it one I purchased.

Summary of Outpost operations: I broke one half of the handle to the output on my trailer’s wastewater pipes because it was frozen. It is still usable. Sergii’s hot water to the bathroom froze last night. It’s been very cold for this early in the season. I’m also using my wall unit to supplement the heat in my house.

Summary of GreenHab operations: Shade cloth was removed. Supplemental light 10-2 pm. Tomatoes all moved to the bed frame. We transplanted the ones from the herb garden into pots as they were not thriving. We removed some of the snow peas to do this, and I replanted snow peas in the herb garden. I thinned the carrots and we replanted microgreens in one of the two raised beds. We also replanted cilantro in clay pot bottoms as they grow better than way. Tomorrow I will clear out the cilantro pot and plant another microgreen.

Summary of ScienceDome operations: Used all week for SWBU. I will be removing the broken microscopes and taking them to WA for replace. One of the teachers who was here for SWBU has donated a new microscope that I will pick up in SLC as soon as possible.

Summary of RAM operations: Added a spark plug gap measuring device and a new ratchet handle to the tools.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, supplies needed and requests: n/a

Copyright © The Mars Society. All rights reserved. | Main Site